OpenStudy (anonymous):

integral of 3/x^4 dx

OpenStudy (lgbasallote):

hmmm 3 can be pulled out and 1/x^4 = x^-4 yes? do you have an idea how to do it now?

OpenStudy (unklerhaukus):

\[\int\frac{3}{x^4}\text d x\] \[=3\int x^{-4} dx\] add one to the index , divide by the new index dont forget to add C

OpenStudy (anonymous):

= - 1/ x^3

OpenStudy (unklerhaukus):

I said dont forget to add C